I want to put a green carpet in a room which is 10ft by 12ft a plum carpet in a room that is 12ft by 14ft and a blue carpet in a room that is 12ft by 16ft. Calculate the amout of each type of carpeting i would need. Suppose the carpet is $24 per sq yard. Calculate the cost of carpeting the rooms. Always round up to the next number.
Answer by josmiceli(19441)

10 x 12 = 120 ft2
12 x 14 = 168 ft2
12 x 16 = 192 ft2
120 + 168 + 192 = 480 ft2
Carpet costs $24 / yd2
(24 / yd2) / (9 ft2 / yd2) = $2.67 / ft2
(480 ft2)(2.67 / ft2) = $1280 answer
another way is
(480 ft2)/ (9 ft2 / yd2) = 53.33 yd2
(53.33 yd2)($24 / yd2) = $1280 answer